# When a Reinforcement Learning Agent Encounters Unknown Unknowns

### v1

@@ 4.4 Learning an Agent with Growing Awareness to Unknown Unknowns
We now design an optimisation method for training our EMDP-GA model for obtaining an agent with growing awareness to handle the unknown unknowns. We adapt upper confidence bound momentum Q-learning (UCBMQ) menard2021ucb to the growing awareness, termed as UCBMQ-GA.
@@
For any transition function $p$ and any function $f:S\rightarrow\mathbb{R}$, we define $pf(s,a)=\mathbb{E}_{s^{\prime}\sim p(\cdot|s,a)}[f(s^{\prime})].$ We denote by $p_{h}^{t}$ the Dirac distribution concentrated at $(s_{h+1}^{t})$, and then, $p_{h}^{t}f(s,a)=f(s_{h+1}^{t})$. The procedure of UCBMQ-GA for training EMDA-GA is also explained below. More details are shown in Appendix C.
@@
$\tilde{Q}^{t-1}_{h}(s,a)=\begin{cases}Q^{t-1}_{h}(s,a),&s\in\mathcal{S}_{t-1},\\ Q^{t-1}_{h,avg}(a),&s\in\mathcal{S}_{t}-\mathcal{S}_{t-1},\end{cases}$ $\ \ \ \tilde{\overline{V}}^{t-1}_{h}(s)=\begin{cases}\overline{V}^{t-1}_{h}(s),&s\in\mathcal{S}_{t-1},\\ \overline{V}_{h,avg}^{t-1},&s\in\mathcal{S}_{t}-\mathcal{S}_{t-1},\end{cases}\text{ and}$
@@
$\beta_{h}^{t}(s,a)=2\sqrt{\frac{\zeta W_{h}^{t}(s,a)}{n_{h}^{t}(s,a)}}+53H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}+\sum_{k=1}^{t}\frac{\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)}{Hlog(T)n_{h}^{t}(s,a)}p_{h}^{k}\left(\tilde{V}_{h,s,a}^{k-1}-\tilde{\overline{V}}_{h+1}^{k-1}\right)(s,a),$
@@ 5.1 Asymptotical Consistency of Learning Unknown Unknowns
We make the following assumption about the transition function.
Assumption 1 (non-zero probability of transition).
For the transition function, We assume that $1-\left(\frac{\delta}{3}\right)^{\frac{1}{H[\sqrt{T}]}}\leq\frac{1}{S}$, and the following holds
$\forall(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S},P(s^{\prime}|s,a)\geq 1-\left(\frac{\delta}{3}\right)^{\frac{1}{H[\sqrt{T}]}}.$
Remark 2.
When $T$ is large enough, $1-\left(\frac{\delta}{3}\right)^{\frac{1}{H[\sqrt{T}]}}\leq\frac{1}{S}$ holds and guarantees that $P(\cdot|s,a)$ can be a distribution over $\mathcal{S}$. Assumption 1 means that the agent can probably reach any state from the current state when taking any action.
@@
Naturally, $t(s)\geq 0$. We then have an upper bound on the aware moment. The proof is detailed in Appendix F.
Theorem 1.
With probability at least $1-\frac{\delta}{3}$, the aware moment is upper-bounded: $t(s)\leq\sqrt{T}$, where $T$ is the number of learning episodes.
Remark 3.
This theorem implies that, with high probability, the agent becomes aware of all states within $\sqrt{T}$ episodes.
This finding is instrumental in deriving regret upper bounds. Then, we obtain the regret bound as follows. A detailed proof is given in Appendix J.
@@
Compared with SOTA, for $H\leq\frac{A}{S}$, the regret bound of our UCBMQ-GA matches UCBMQ ($\mathcal{O}(\sqrt{SAH^{3}T}+SAH^{4})$) menard2021ucb . For $T\geq H^{5}SA$ and $H\leq\frac{A}{S}$, the regret bound of UCBMQ-GA matches monotonic value propagation (MVP, $\mathcal{O}(\sqrt{H^{3}SAT})$) zhang2024settling . Other methods, upper Bayes-confidence bound value iteration (Bayes-UCBVI) tiapkin2022dirichlet and upper confidence bound-advantage(UCB-Adv) zhang2020almost , are worse than the three methods in either case. Table 1 shows the comparisons.
@@
UCBMQ menard2021ucb $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA)$ $\mathcal{O}(H(S+A)T)$ $\mathcal{O}(HS^{2}A)$
MVP zhang2024settling $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT})$ $\mathcal{O}(HSAT)$ $\mathcal{O}(HS^{2}A)$
UCB-Adv zhang2020almost $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{\frac{33}{4}}S^{2}A^{\frac{3}{2}}T^{\frac{1}{4}})$ $\mathcal{O}(HAT)$ $\mathcal{O}(HSA)$
Bayes-UCBVI tiapkin2022dirichlet $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{3}S^{2}A)$ $\mathcal{O}(BHS^{2}AT)$ $\mathcal{O}(HS^{2}A)$
UCBMQ-GA (ours) $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA+H^{2}S\sqrt{T})$ $\mathcal{O}(H(S+A)T)$ $\mathcal{O}(HS^{2}A)$
@@
The proof is in five steps. $\mathring{\overline{Q}}_{h}^{t}(s,a)$ and $\mathring{\overline{V}}_{h}^{t}(s)$ are defined in Appendix A.
@@
$\hat{R}_{h}^{T}(s,a)\triangleq\sum_{t=0}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)=\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a).$
@@
$\hat{R}_{h}^{T}(s,a)\leq$ $63log(T)\sqrt{\zeta\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}+1754H^{3}log(T)^{2}\zeta$
@@ 5.2 Computational Complexity and Space Complexity
The proof is given in Appendix G. Here, we present an intuitive sketch. The updates of $\overline{Q}_{h}^{t}$, $\overline{V}_{h}^{t}$ and $V_{h,s,a}^{t}$ are carried out in an online manner. At time step $h$ of the episode $t$, if $(s_{h}^{t},a_{h}^{t})\neq(s,a)$, both learning rate $\alpha_{H}^{t}(s,a)$ and momentum term $\gamma_{h}^{t}(s,a)$ become zero. Thus, we do not need to update the functions on such $(s,a)$ pairs. Furthermore, the expansions of $Q^{t-1}_{h}$, $\overline{V}^{t-1}_{h}$, and $V^{t-1}_{h,s,a}$ in each episode depend on their corresponding averaged values on $\mathcal{S}_{t-1}$.
@@
A detailed proof is presented in Appendix G. Intuitively, this lemma holds because we need to store all the bias-value function $V_{h,s,a}^{t}(s^{\prime})$.
@@ Appendix A Additional Notations and Preliminaries
Dirac distribution concentrated at $(s_{h+1}^{t})$ $p_{h}^{t}(s^{\prime}|s,a),\ s,s^{\prime}\in\mathcal{S}_{t}$ $\mathring{p}_{h}^{t}(s^{\prime}|s,a),\ s,s^{\prime}\in\mathcal{S}$
Estimated Q-value function $Q_{h}^{t}(s,a),\ s\in\mathcal{S}_{t},\ a\in\mathcal{A}$ $\tilde{Q}_{h}^{t}(s,a),\ s\in\mathcal{S}_{t+1},\ a\in\mathcal{A}$ $\mathring{Q}_{h}^{t}(s,a),\ s\in\mathcal{S},\ a\in\mathcal{A}$
Bias-value function $V_{h,s,a}^{t}(s^{\prime}),\ s,s^{\prime}\in\mathcal{S}_{t}$ $\tilde{V}_{h,s,a}^{t}(s^{\prime}),\ s,s^{\prime}\in\mathcal{S}_{t+1}$ $\mathring{V}_{h,s,a}^{t}(s^{\prime}),\ s,s^{\prime}\in\mathcal{S}$
Upper bound on Q-value function $\overline{Q}_{h}^{t}(s,a),\ s\in\mathcal{S}_{t},\ a\in\mathcal{A}$ $\tilde{\overline{Q}}_{h}^{t}(s,a),\ s\in\mathcal{S}_{t+1},\ a\in\mathcal{A}$ $\mathring{\overline{Q}}_{h}^{t}(s,a),\ s\in\mathcal{S},\ a\in\mathcal{A}$
Upper bound on value function $\overline{V}_{h}^{t}(s),\ s\in\mathcal{S}_{t}$ $\tilde{\overline{V}}_{h}^{t}(s),\ s\in\mathcal{S}_{t+1}$ $\mathring{\overline{V}}_{h}^{t}(s),\ s\in\mathcal{S}$
@@
$n_{h}^{t}(s,a)Q_{h}^{t}(s,a)=$ $\chi_{h}^{t}(s,a)[r_{h}(s,a)+p_{h}^{t}\tilde{\overline{V}}_{h+1}^{t-1}(s,a)+\mathring{\gamma}_{h}^{t}(s,a)p_{h}^{t}(\tilde{\overline{V}}_{h+1}^{t-1}-\tilde{V}_{h,s,a}^{t-1})(s,a)]$
@@
$n_{h}^{t}(s,a)Q_{h}^{t}(s,a)=$ $\chi_{h}^{t}(s,a)[r_{h}(s,a)+p_{h}^{t}\tilde{\overline{V}}_{h+1}^{t-1}(s,a)+\mathring{\gamma}_{h}^{t}(s,a)p_{h}^{t}(\tilde{\overline{V}}_{h+1}^{t-1}-\tilde{V}_{h,s,a}^{t-1})(s,a)]$
@@
$Q_{h}^{t}(s,a)$ $=r_{h}(s,a)+\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[p_{h}^{k}\tilde{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)p_{h}^{k}(\tilde{\overline{V}}_{h+1}^{k-1}-\tilde{V}_{h,s,a}^{k-1})(s,a)].$
@@
$Q_{h}^{t}(s,a)=r_{h}(s,a)+\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a)+\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)].$
@@
$=\eta_{h}^{t}(s,a)\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}\overline{V}_{h+1}^{t-1}(\hat{s})+(1-\eta_{h}^{t}(s,a))\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}V_{h,s,a}^{t-1}(\hat{s})$
$\geq\eta_{h}^{t}(s,a)\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}\overline{V}_{h+1}^{t-1}(\hat{s})+(1-\eta_{h}^{t}(s,a))\frac{1}{|\mathcal{S}_{t-1}|}\sum_{\hat{s}\in\mathcal{S}_{t-1}}\overline{V}_{h+1}^{t-1}(\hat{s})$
@@ Appendix B A Brief Introduction to Upper Confidence Bound Momentum Q-learning
$\beta_{h}^{t}(s,a)=2\sqrt{\frac{\zeta W_{h}^{t}(s,a)}{n_{h}^{t}(s,a)}}+53H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}+\sum_{k=1}^{t}\frac{\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)}{Hlog(T)n_{h}^{t}(s,a)}p_{h}^{k}(V_{h,s,a}^{k-1}-\overline{V}_{h+1}^{k-1})(s,a).$

### v2

@@ 4.4 Learning an Agent with Growing Awareness to Unknown Unknowns
We now design an optimisation method for training our EMDP-GA model for obtaining an agent with growing awareness to handle the unknown unknowns. We adapt upper confidence bound momentum Q-learning (UCBMQ) [24] to the growing awareness, termed as UCBMQ-GA.
@@
For any transition function $p$ and any function $f:S\rightarrow\mathbb{R}$, we define $pf(s,a)=\mathbb{E}_{s^{\prime}\sim p(\cdot|s,a)}[f(s^{\prime})].$ We denote by $p_{h}^{t}$ the Dirac distribution concentrated at $(s_{h+1}^{t})$, and then, $p_{h}^{t}f(s,a)=f(s_{h+1}^{t})$. The procedure of UCBMQ-GA for training EMDA-GA is also explained below. More details are shown in C.
@@
$\text{and }\tilde{V}^{t-1}_{h,s,a}(s^{\prime})=\begin{cases}V^{t-1}_{h,s,a}(s^{\prime}),&s,s^{\prime}\in\mathcal{S}_{t-1},\\ V_{h,s,a,avg}^{t-1},&s\in\mathcal{S}_{t-1},s^{\prime}\in\mathcal{S}_{t}-\mathcal{S}_{t-1},\\ V_{h,a,avg}^{t-1}(s^{\prime}),&s\in\mathcal{S}_{t}-\mathcal{S}_{t-1},s^{\prime}\in\mathcal{S}_{t-1},\\ V_{h,a,avg}^{t-1},&s,s^{\prime}\in\mathcal{S}_{t}-\mathcal{S}_{t-1}.\end{cases}$
@@
$\beta_{h}^{t}(s,a)=$ $2\sqrt{\frac{\zeta W_{h}^{t}(s,a)}{n_{h}^{t}(s,a)}}+53H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
$+\sum_{k=1}^{t}\frac{\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)}{Hlog(T)n_{h}^{t}(s,a)}p_{h}^{k}\left(\tilde{V}_{h,s,a}^{k-1}-\tilde{\overline{V}}_{h+1}^{k-1}\right)(s,a),$
@@ 5.1 Asymptotical Consistency of Learning Unknown Unknowns
Naturally, $t(s)\geq 0$. We then have an upper bound on the aware moment. The proof is detailed in LABEL:proof_of_first_existence_lemma.
This finding is instrumental in deriving regret upper bounds. Then, we obtain the regret bound as follows. A detailed proof is given in I.
@@
Compared with SOTA, for $H\leq\frac{A}{S}$, the regret bound of our UCBMQ-GA matches UCBMQ ($\mathcal{O}(\sqrt{SAH^{3}T}+SAH^{4})$) [24]. For $T\geq H^{5}SA$ and $H\leq\frac{A}{S}$, the regret bound of UCBMQ-GA matches monotonic value propagation (MVP, $\mathcal{O}(\sqrt{H^{3}SAT})$) [38]. Other methods, upper Bayes-confidence bound value iteration (Bayes-UCBVI) [31] and upper confidence bound-advantage(UCB-Adv) [40], are worse than the three methods in either case. Table 1 shows the comparisons.
@@
UCBMQ [24] $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA)$ $\mathcal{O}(H(S+A)T)$ $\mathcal{O}(HS^{2}A)$
MVP [38] $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT})$ $\mathcal{O}(HSAT)$ $\mathcal{O}(HS^{2}A)$
UCB-Adv [40] $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}$ $H^{\frac{33}{4}}S^{2}A^{\frac{3}{2}}T^{\frac{1}{4}})$ $\mathcal{O}(HAT)$ $\mathcal{O}(HSA)$
Bayes-UCBVI [31] $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{3}S^{2}A)$ $\mathcal{O}(BHS^{2}AT)$ $\mathcal{O}(HS^{2}A)$
UCBMQ-GA (ours) $\tilde{\mathcal{O}}(\sqrt{H^{3}SAT}+H^{4}SA$ $H^{2}S\sqrt{T})$ $\mathcal{O}(H(S+A)T)$ $\mathcal{O}(HS^{2}A)$
@@
The proof is in five steps. $\mathring{\overline{Q}}_{h}^{t}(s,a)$ and $\mathring{\overline{V}}_{h}^{t}(s)$ are defined in A.
@@
$\hat{R}_{h}^{T}(s,a)$ $\triangleq\sum_{t=0}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a)$
$=\sum_{t=t(s)-1}^{T-1}\chi_{h}^{t+1}(s,a)(\mathring{\overline{Q}}_{h}^{t}-Q_{h}^{\pi^{t+1}})(s,a).$
@@
$\hat{R}_{h}^{T}(s,a)\leq$ $63log(T)\sqrt{\zeta\sum_{t=t(s)-1}^{T-1}\overline{p}_{h}^{t+1}(s,a)Var_{p_{h}}(V_{h+1}^{\pi^{t+1}})(s,a)}$
$+1754H^{3}log(T)^{2}\zeta$
@@ 5.2 Computational Complexity and Space Complexity
The proof is given in F. Here, we present an intuitive sketch. The updates of $\overline{Q}_{h}^{t}$, $\overline{V}_{h}^{t}$ and $V_{h,s,a}^{t}$ are carried out in an online manner. At time step $h$ of the episode $t$, if $(s_{h}^{t},a_{h}^{t})\neq(s,a)$, both learning rate $\alpha_{H}^{t}(s,a)$ and momentum term $\gamma_{h}^{t}(s,a)$ become zero. Thus, we do not need to update the functions on such $(s,a)$ pairs. Furthermore, the expansions of $Q^{t-1}_{h}$, $\overline{V}^{t-1}_{h}$, and $V^{t-1}_{h,s,a}$ in each episode depend on their corresponding averaged values on $\mathcal{S}_{t-1}$.
@@
A detailed proof is presented in F. Intuitively, this lemma holds because we need to store all the bias-value function $V_{h,s,a}^{t}(s^{\prime})$.
@@ Appendix A Additional Notations and Preliminaries
Dirac distribution concentrated at $(s_{h+1}^{t})$
$p_{h}^{t}(s^{\prime}|s,a),\ s,s^{\prime}\in\mathcal{S}_{t}$
$\mathring{p}_{h}^{t}(s^{\prime}|s,a),\ s,s^{\prime}\in\mathcal{S}$
Estimated Q-value function
$Q_{h}^{t}(s,a),\ s\in\mathcal{S}_{t},\ a\in\mathcal{A}$
$\tilde{Q}_{h}^{t}(s,a),\ s\in\mathcal{S}_{t+1},\ a\in\mathcal{A}$
$\mathring{Q}_{h}^{t}(s,a),\ s\in\mathcal{S},\ a\in\mathcal{A}$
Bias-value function
$V_{h,s,a}^{t}(s^{\prime}),\ s,s^{\prime}\in\mathcal{S}_{t}$
$\tilde{V}_{h,s,a}^{t}(s^{\prime}),\ s,s^{\prime}\in\mathcal{S}_{t+1}$
$\mathring{V}_{h,s,a}^{t}(s^{\prime}),\ s,s^{\prime}\in\mathcal{S}$
Upper bound on Q-value function
$\overline{Q}_{h}^{t}(s,a),\ s\in\mathcal{S}_{t},\ a\in\mathcal{A}$
$\tilde{\overline{Q}}_{h}^{t}(s,a),\ s\in\mathcal{S}_{t+1},\ a\in\mathcal{A}$
$\mathring{\overline{Q}}_{h}^{t}(s,a),\ s\in\mathcal{S},\ a\in\mathcal{A}$
Upper bound on value function
$\overline{V}_{h}^{t}(s),\ s\in\mathcal{S}_{t}$
$\tilde{\overline{V}}_{h}^{t}(s),\ s\in\mathcal{S}_{t+1}$
$\mathring{\overline{V}}_{h}^{t}(s),\ s\in\mathcal{S}$
@@
$\ \ \ \ n_{h}^{t}(s,a)Q_{h}^{t}(s,a)$
$=\chi_{h}^{t}(s,a)[r_{h}(s,a)+p_{h}^{t}\tilde{\overline{V}}_{h+1}^{t-1}(s,a)+\mathring{\gamma}_{h}^{t}(s,a)p_{h}^{t}(\tilde{\overline{V}}_{h+1}^{t-1}-\tilde{V}_{h,s,a}^{t-1})(s,a)]$
@@
$\ \ \ \ n_{h}^{t}(s,a)Q_{h}^{t}(s,a)$
$=\chi_{h}^{t}(s,a)[r_{h}(s,a)+p_{h}^{t}\tilde{\overline{V}}_{h+1}^{t-1}(s,a)+\mathring{\gamma}_{h}^{t}(s,a)p_{h}^{t}(\tilde{\overline{V}}_{h+1}^{t-1}-\tilde{V}_{h,s,a}^{t-1})(s,a)]$
@@
$Q_{h}^{t}(s,a)=$ $r_{h}(s,a)+\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[p_{h}^{k}\tilde{\overline{V}}_{h+1}^{k-1}(s,a)+$
$\mathring{\gamma}_{h}^{k}(s,a)p_{h}^{k}(\tilde{\overline{V}}_{h+1}^{k-1}-\tilde{V}_{h,s,a}^{k-1})(s,a)].$
@@
$Q_{h}^{t}(s,a)=$ $r_{h}(s,a)+\frac{1}{n_{h}^{t}(s,a)}\sum_{k=t(s)}^{t}\chi_{h}^{k}(s,a)[\mathring{p}_{h}^{k}\mathring{\overline{V}}_{h+1}^{k-1}(s,a)$
$+\mathring{\gamma}_{h}^{k}(s,a)\mathring{p}_{h}^{k}(\mathring{\overline{V}}_{h+1}^{k-1}-\mathring{V}_{h,s,a}^{k-1})(s,a)].$
@@
[…]
@@ Appendix B A Brief Introduction to Upper Confidence Bound Momentum Q-learning
$\beta_{h}^{t}(s,a)=$ $2\sqrt{\frac{\zeta W_{h}^{t}(s,a)}{n_{h}^{t}(s,a)}}+53H^{3}\frac{\zeta log(T)}{n_{h}^{t}(s,a)}$
$+\sum_{k=1}^{t}\frac{\chi_{h}^{k}(s,a)\mathring{\gamma}_{h}^{k}(s,a)}{Hlog(T)n_{h}^{t}(s,a)}p_{h}^{k}(V_{h,s,a}^{k-1}-\overline{V}_{h+1}^{k-1})(s,a).$